OpenStudy (anonymous):

integral sin(1/x) dx ?

OpenStudy (anonymous):

xsin(1/x) - Ci(1/x) + C

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

you can't integrate this thing. where did the question come from?

OpenStudy (anonymous):

\[\int\limits 8x^2 \sin (1/10 x) dx\]

OpenStudy (anonymous):

thats where the question came from

OpenStudy (anonymous):

for the love of pete, wolfram doesn't give an answer! it gives the integral in terms of another integral.

OpenStudy (anonymous):

\[8\int x^2\sin(\frac{1}{10}x)dx\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it is really \[\sin(\frac{x}{10})\] not \[\sin(\frac{1}{10x})\]

OpenStudy (anonymous):

hahahahahah sorry

OpenStudy (anonymous):

integrate by parts twice. pain in the butt, but if you want to use wolfram and follow the steps at least you will get a real answer this time

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

here you can look at this. but the basic idea is to integrate sin and differen tiate x^2. then you will have to do it once more http://www.wolframalpha.com/input/?i=integral+x^2sin%28x%2F%2810%29%29